1. Firm A just purchased a machine for $ 1,517,824 . If the salvage value of this machine is $ 352,536 and if the estimated life of the machine is 14 years, what is the annual depreciation on this machine? (Round your answer to the nearest whole dollar and record without a dollar sign or a comma. For example, record $34,253.85431 as 34254).

2. Firm X purchased a piece of equipment exactly 9 years ago. The piece of equipment had a purchase price of $ 8,985,999 , a salvage value of $ 344,768 , and an estimated useful life of 22 years. What is the current book value of this piece of equipment owned by Firm X? (Round your answer to the nearest whole dollar and record without a dollar sign or a comma. For example, record $34,253.85431 as 34254).
Answer #1

1.Annual depreciation=(Cost-Salvage value)/Useful Life

=(1,517,824-352,536)/14

=83235(Approx).

2.Annual depreciation=(Cost-Salvage value)/Useful Life

=(8,985,999-344,768)/22

=$392,783.23(Approx)

Hence current book value=Cost-Accumulated depreciation

=8,985,999-(392,783.23*9)

=5450950(Approx).